### v1

@@ Abstract
The trend of employing training-free methods for point cloud recognition is becoming increasingly popular due to its significant reduction in computational resources and time costs. However, existing approaches are limited as they typically extract either geometric or semantic features. To address this limitation, we propose a novel method that integrates both geometric and semantic features, thereby enhancing the comprehensive understanding of point clouds. For the geometric branch, we adopt a non-parametric strategy to extract geometric features. In the semantic branch, we leverage a model pre-trained through contrastive learning and aligned with text features to obtain semantic features. Experimental results demonstrate that our method outperforms existing state-of-the-art training-free approaches on several popular benchmark datasets, including ModelNet and ScanObiectNN.
@@
Point Cloud, Training-free, Feature Fusion, Few-shot Learning
@@ III-C Geometric Feature Enhancement
These enhanced features strengthen the spatial information in our model, improving performance in classification tasks.
@@ IV-B Comparative Experiments
We compare our approach with existing training-free methods, as shown in Table I. Our approach demonstrate outstanding performance across the three benchmark datasets: ModelNet10[25], ModelNet40[25], and ScanObjectNN[26]. We achieve state-of-the-art (SOTA) performance in both full-shot and 16-shot settings. This indicates that our approach is not only effective on synthetic data but also maintains high accuracy and stability in challenging real-world scenarios. Furthermore, the experimental results show that our approach can achieve efficient classification results in a very short time, relying solely on feature memory without the need for additional training. This further validates the effectiveness and practicality of our architecture.
@@ V Conclusion
This work presents a novel approach to point cloud recognition that integrates both geometric and semantic features, addressing the limitations of existing training-free methods. Our approach outperforms state-of-the-art methods on benchmark datasets like ModelNet and ScanObjectNN, demonstrating the effectiveness of combining these two feature types for comprehensive point cloud analysis.

### v2

@@ Abstract
The trend of employing training-free methods for point cloud recognition is becoming increasingly popular due to its significant reduction in computational resources and time costs. However, existing approaches are limited as they typically extract either geometric or semantic features. To address this limitation, we are the first to propose a novel training-free method that integrates both geometric and semantic features. For the geometric branch, we adopt a non-parametric strategy to extract geometric features. In the semantic branch, we leverage a model aligned with text features to obtain semantic features. Additionally, we introduce the GFE module to complement the geometric information of point clouds and the MFF module to improve performance in few-shot settings. Experimental results demonstrate that our method outperforms existing state-of-the-art training-free approaches on mainstream benchmark datasets, including ModelNet and ScanObiectNN.
@@
Point Cloud, Training-free, Feature Fusion, Few-shot Classification
@@ III-C Geometric Feature Enhancement
These enhanced features provide richer spatial information for our model, improving performance in classification tasks.
@@ IV-B Comparative Experiments
We compare our approach with existing training-free and fine-tuned methods. As shown in Table I, our approach shows state-of-the-art performance in both full-shot and 16-shot settings on three benchmark datasets: ModelNet10[25], ModelNet40[25], and ScanObjectNN[26]. This indicates that our approach is not only effective on synthetic data but also maintains high accuracy and stability in challenging real-world scenarios. Furthermore, our approach can achieve accurate classification results under limited resources, relying solely on feature memory without training.
@@ V Conclusion
This work, for the first time, presents a novel approach to point cloud recognition that integrates both geometric and semantic information in a training-free framework. The efficiency and understanding of point clouds are improved by selecting key features from the feature memory and enriching geometric feature extraction. Extensive evaluations on benchmark datasets such as ModelNet and ScanObjectNN demonstrate that our method achieves state-of-the-art performance, showing the effectiveness of our feature fusion strategy for accurate point cloud recognition without training.